# Efficient Large Language Models: A Survey

### v2

@@ 2.1.2. Parameter Pruning
Parameter pruning compresses LLMs by removing redundant model weights. Parameter pruning methods for LLMs can be categorized into structured pruning and unstructured pruning.
@@
Structured Pruning. Structured pruning focuses on pruning structured patterns such as groups of consecutive parameters or hierarchical structures such as rows, columns, or sub-blocks of the LLM weight matrices. For example, LLM-Pruner (Ma et al., 2023) introduces a task-agnostic structured pruning strategy that selectively eliminates non-essential interconnected structures using gradient information. It utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for LLaMA (Touvron et al., 2023a), and uses LoRA (Hu et al., 2022) to recover performance after pruning, showing competitive zero shot performance. Sheared LLaMA (Xia et al., 2023) proposes two techniques. The first technique is targeted structured pruning, which prunes a larger model to a designated target shape by eliminating layers, heads, and intermediate and hidden dimensions in a end-to-end fashion. The second technique is dynamic batch loading, which dynamically alters the components of the sampled data in each training batch based on losses in various domains. Through these two techniques, Sheared LLaMA is able to prune the LLaMA2-7B model down to 1.3B parameters. LoRAPrune (Zhang et al., 2023c) introduces a LoRA-based pruning criterion using LoRA’s weights and gradients instead of pre-trained weights’ gradients for importance estimation. By employing a structured iterative pruning process to eliminate excess channels and heads, LoRAPrune outperforms LLM-Pruner in efficiency at a 50% compression rate.
@@
Unstructured Pruning. Unstructured pruning, on the other hand, focuses on pruning model weights individually, and thus has much more flexibility compared to structured pruning. Frantar and Alistarh (2023) present SparseGPT, a one-shot LLM pruning approach that does not require retraining. It formulates pruning as a sparse regression problem and solves it by utilizing an approximate solver based on the inversion of the Hessian matrix. In doing so, SparseGPT reaches 60 unstructured sparsity even on models such as OPT-135B while experiencing only a slight reduction in perplexity. Sun et al. (2023) propose Wanda which prunes weights based on the product values of weight magnitudes and their respective input activations. Compared to SparseGPT, Wanda neither relies on second-order information nor necessitates weight update, and performs competitively against SparseGPT. Shao et al. (2023) propose to utilize Hessian sensitivity-aware mixed sparsity pruning to achieve a minimum of 50% sparsity in LLMs without retraining. This method adaptively assigns sparsity based on sensitivity to minimize the error induced by pruning while preserving the overall level of sparsity.
2.1.3. Low-Rank Approximation
Low-rank approximation compresses LLMs by approximating the weight matrix of LLMs with low-rank matrices and such that , where , , and is typically much smaller than . In doing so, low-rank approximation reduces the number of parameters and enhances computational efficiency. In particular, Xu et al. (2023c) introduce TensorGPT which compresses the embedding layers of LLMs using Tensor-Train Decomposition (TTD). It transforms and breaks down each token embedding and creates an efficient embedding format named Matrix Product State (MPS) that can be efficiently computed in a distributed manner. LoSparse (Li et al., 2023g) aims to compress the coherent and expressive components within neurons through low-rank approximation while eliminating the incoherent and non-expressive elements via pruning the sparse matrix. It uses iteration training to calculate the important score of column neurons for pruning, outperforming conventional iterative pruning methods.
2.1.4. Knowledge Distillation
Knowledge Distillation (KD) compresses LLMs by training a smaller student model to emulate the performance of the LLM as the teacher model such that the student model is computationally less expansive yet maintains a high level of performance similar to the teacher model. KD for LLMs can be categorized into white-box KD methods and black-box KD methods.
@@ 2.1.4. Knowledge Distillation
White-Box Knowledge Distillation. White-box KD refers to KD techniques where the parameters or logits of the teacher LLM are used in the distillation process (Gou et al., 2021). For example, Baby LLaMA (Timiryasov and Tastet, 2023) trains an ensemble of GPT-2 and a collection of smaller LLaMA-1 models using the BabyLM dataset of 10M words. This ensemble is then distilled into a compact LLaMA model with 58 million parameters, which outperforms both its original teacher models as well as a comparable model that was trained without the use of distillation. Gu et al. (2023) observe that conventional KD objectives, such as Kullback-Leibler divergence (KLD), may not be well suited for open text generation tasks due to their more complex output spaces compared to classification tasks. To address this issue, they propose MiniLLM that minimizes reverse KLD using the gradient of the objective function through policy gradient techniques (Sutton et al., 1999). This approach surpasses the performance of standard KD benchmarks on the 13-billion-parameter LLaMA-1 model (Touvron et al., 2023a). Similarly, generalized knowledge distillation (GKD) (Agarwal et al., 2023) addresses the issue of distribution mismatch by drawing output sequences from the student model during training. GKD tackles the problem of model under-specification by optimizing different divergence measures, like reverse KL. This approach aims to produce samples from the student model that are probable within the teacher model’s distribution. KPTD (Padmanabhan et al., 2023) demonstrates that KD methods can successfully transfer and disseminate knowledge from entity definitions into the parameters of a pre-trained language model. Specifically, it creates a transfer set by prompting the language model to generate text based on the definition of the entity. Then the models’ parameters are updated to align the distribution of the student language model with that of the teacher model. TED (Liang et al., 2023) introduces a technique for layer-specific task distillation. It uses specially designed filters to align the internal states of both student and teacher models in each layer. These filters extract the relevant knowledge from the internal states that is beneficial for the specific task. TED shows considerable and steady gains in performance on both continual pre-training and fine-tuning. TSLD (Kim et al., 2023c) leverages token-level distillation to enhance QAT, which overcomes the limitations of layer-to-layer KD in token prediction recovery by reforming intermediate representation and has successfully applied QAT to LLMs. Lastly, MiniMA (Zhang et al., 2023e) proposes a viewport towards the capacity gap in distilling LLMs, converting it into a principle through analysis and introducing a 3B Language Model that sets a new benchmark for compute-performance pareto frontier.
Black-Box Knowledge Distillation. Different from white-box KD, in black-box KD, only the outputs generated from the teacher LLM are used in the distillation process. Inspired by MetaICL and MetalICL (Chen et al., 2022b; Min et al., 2021), where the language model is meta-trained in a wide range of tasks using in-context learning objectives and then fine-tuned for unseen tasks through in-context learning, Multitask-ICT (Huang et al., 2022) introduces a concept known as in-context learning distillation. This method aims to transfer the few-shot learning capabilities from the LLM teacher to the student model. Similarly, LI et al. (2022) introduce a hybrid prompting technique that employs multi-task learning along with explanations generated by GPT-3 text-davinci-002 version (OpenAI, 2023). This method is used to distill explanations into smaller models, achieving consistent and significant improvements over strong single-task fine-tuning benchmarks in different scenarios. Lion (Jiang et al., 2023a) introduces an adversarial distillation architecture aimed at enhancing the efficiency of knowledge transfer by incrementally improving the skill level of the student model. Specifically, it prompts LLMs to recognize challenging instructions and create new complex instructions for the student model, thereby establishing a three-phase adversarial cycle involving imitation, discrimination, and generation. DISCO (Chen et al., 2023b) involves prompting a general LLM to produce phrasal perturbations. These generated perturbations are then filtered by a specialized teacher model to distill high-quality counterfactual data into smaller student models, allowing the smaller models to learn causal representations more reliably. Recently, some studies have shown that chain-of-thought (CoT) prompting can elicit language models to solve complex reasoning tasks step by step, with the aim of transfer this ability from LLMs into smaller models through black-box KD. For example, Fu et al. (2023c) aim to enhance the CoT math reasoning capabilities of smaller models. Specifically, they employ a method that involves instruct-tuning an student model (FlanT5) by distilling the reasoning pathways found in the GSM8K dataset from a LLM teacher (GPT-3.5 code-davinci-002 (Chen et al., 2021b)). The small model is then selected based on its average performance on three separate, withheld math reasoning datasets to confirm its ability to generalize well to new, out-of-distribution scenarios. Likewise, Distilling Step-by-Step (Hsieh et al., 2023) claims that to match the performance of LLMs, fine-tuning and distilling smaller models require substantial amounts of training data. To address this, it proposes a technique that uses CoT prompting to extract LLM rationales for extra guidance in training smaller models within a multi-task setting, achieving better performance compared to few shot prompted LLMs. Fine-tune-CoT (Ho et al., 2023) utilizes existing zero-shot CoT prompting techniques (Kojima et al., 2023) to create rationales from LLMs. These rationales are then used to fine-tune smaller student models. The approach also introduces diverse reasoning, a method that employs stochastic sampling to generate a variety of reasoning solutions from teacher models, which serves to enrich the training data for the student models. SOCRATIC CoT (Shridhar et al., 2022) employs a method that breaks down the original problem into a series of smaller tasks and utilizes this decomposition to direct the intermediate steps of reasoning. This approach is used to train a pair of smaller, distilled models: one that specializes in dissecting the problem and another focused on solving these sub-problems. SCOTT (Wang et al., 2023f) uses rationales generated by LLMs to train a student model under a counterfactual reasoning framework. This approach ensures that the student model does not overlook the provided rationales, thereby preventing it from making inconsistent predictions. SCoTD (Li et al., 2023a) presents a method called symbolic CoT distillation. It involves drawing CoT rationales from a LLM using unlabeled data instances. A smaller model is then trained to predict both the sampled rationales and the associated labels. Lastly, Peng et al. (2023b) utilize GPT-4 as a teacher model to generate English and Chinese instruction-based datasets to refine student LLMs such as LLaMA. Their results show that the 52K data points generated by GPT-4 are able to improve zero-shot performance compared to instruction-following data generated from previous state-of-the-art models.
@@
| OPT (Zhang et al., 2022d) | 175B | 180B tokens | 992 A100-80G | - |
| BLOOM (Scao et al., 2022) | 176B | 366B tokens | 384 A100-80G | 105 days |
| GLM (Zeng et al., 2022) | 130B | 400B tokens | 786 A100-40G | 60 days |
@@
| Gopher (Rae et al., 2021) | 280B | 300B tokens | 1024 A100 | 13.4 days |
@@
| WeLM (Su et al., 2022b) | 10B | 300B tokens | 128 A100-40G | 24 days |
@@
2.2. Efficient Pre-Training
As shown in Table 1, pre-training LLMs incurs high costs. Efficient pre-training aims to enhance the efficiency and reduce the cost of the LLM pre-training process. As summarized in Figure 7, efficient pre-training techniques can be grouped into four categories: mixed precision acceleration, scaling models, initialization techniques, and optimization strategies.
@@ 2.2. Efficient Pre-Training
Mixed Precision Acceleration, fill=blue!10
Automatic Mixed Precision (AMP) (Micikevicius
et al., 2017; Facebook AI
Research (FAIR), 2023; Rae
et al., 2021),
Brain Floating Point (BF16) (Kalamkar
et al., 2019; Burgess et al., 2019),
GACT (Liu et al., 2022d),
Mesa (Pan
et al., 2021), leaf, text width=48em
@@
Gong
et al. (2019), MSLT(Yang
et al., 2020), CompoundGrow(Gu
et al., 2021), Knowledge Inheritance (Qin et al., 2021), Staged Training (Shen et al., 2022),
@@
Mango (Pan
et al., 2023),
Yao
et al. (2023d),
Growth Strategy (Li et al., 2023d), leaf, text width=45em
@@
Kumar (2017), Fixup (Zhang
et al., 2019), ZerO (Zhao
et al., 2021), SkipInit (De and Smith, 2020),
ReZero (Bachlechner et al., 2021), T-Fixup (Huang
et al., 2020), DeepNet (Wang
et al., 2022c), leaf, text width=23em
Optimization Strategies, fill=blue!10
Lion (Chen et al., 2023f),
Sophia (Liu
et al., 2023d), leaf, text width=10em
@@
ZeRO-Infinity (Rajbhandari et al., 2021a), leaf, text width=28em
@@
Mixed Precision Acceleration. Mixed precision acceleration enhances pre-training efficiency by using the low-precision model for forward and backward propagation and converting the calculated low-precision gradients to high-precision ones for updating the original high-precision weights. For example, Micikevicius et al. (2017) propose Automatic Mixed Precision (AMP) to keep a master copy of weights in full-precision FP32 for updates, whereas weights, activations, and gradients are stored in FP16 for arithmetic operations. Notably, the improved version of AMP (Facebook AI Research (FAIR), 2023) optimizer has eliminated the copy of FP32 weights, but the optimizer (AdamW) still use FP32 internally. However, Rae et al. (2021) demonstrate that FP16 results in accuracy loss. To counteract this performance drop, Brain Floating Point (BF16) was proposed (Kalamkar et al., 2019; Burgess et al., 2019), which achieves better performance by assigning more bits to the exponent and fewer to the significant bits. Lastly, recent studies (Pan et al., 2021; Liu et al., 2022d) have shown that combining mixed-precision acceleration with activation compressed training (ACT) can further facilitate memory-efficient Transformer pre-training.
@@
Scaling Models. Techniques based on scaling models accelerate pre-training convergence and reduce training costs by using the weights of a small model to scale up to a large model. For example, Gong et al. (2019) introduce Progressive Stacking to transfer knowledge from a simpler model to a more complex one and then uses progressive stacking to enhance the model’s training efficiency and convergence speed. Yang et al. (2020) observe that as the depth of the model increases through progressive stacking, the training speed however decreases. To address this issue, they propose multi-stage layer training (MSLT), which only updates the output and newly introduced top encoder layers while keeping the previously trained layers unchanged. Once all the layers have been trained, MSLT fine-tunes the entire model by updating each layer in just 20% of the total steps, making it more time-efficient than the traditional progressive stacking approach. Gu et al. (2021) introduce CompoundGrow, which begins with the training of a small model and incrementally expands it using a mix of model growth techniques, including increasing input length, model breadth, and depth, leading to an acceleration in the pre-training process by up to 82.2%. Qin et al. (2021) propose Knowledge Inheritance which employs knowledge distillation as an auxiliary supervision during pre-training. This aids in effectively training a larger model from a smaller teacher model, thereby enhancing both the speed of pre-training and the generalization ability. Shen et al. (2022) introduce Staged Training that begins with a small model and progressively increases its depth and breadth through a growth operator, which includes model parameters, the state of the optimizer, and the learning rate schedule. By starting each phase with the results from the previous one, it effectively reuses computation, leading to a more efficient training process. Chen et al. (2021c) propose function-reserving initialization (FPI) and advanced knowledge initialization (AKI) to transfer the knowledge of a smaller pre-trained model to a large model to improve the pre-training efficiency of the large model. Specifically, FPI gives the larger model a behavior similar to that of the smaller model, laying a strong basis for optimization; and AKI promotes faster convergence by replicating weights from higher layers. Wang et al. (2023d) propose Linear Growth Operator (LiGO) that linearly maps the parameters of a smaller model to initiate a larger one, using a composition of width-and depth-growth operators, further enhanced with Kronecker factorization to capture architectural knowledge. Mango (Pan et al., 2023) introduces a technique that establishes a linear relationship between each weight of the target model and all weights of the pretrained model to boost acceleration capabilities. It also employs multi-linear operators to decrease computational and spatial complexity during pre-training. Drawing from these scaling techniques and the progressive pre-training (Yao et al., 2023d), recent LLMs like FLM-101B (Li et al., 2023d) introduce a growth strategy to cut LLM training costs by expanding model structures offline and resuming from the previous stage’s smaller model checkpoint.
@@
Initialization Techniques. Initialization plays a key role in enhancing the efficiency of LLM pre-training since a good initialization can accelerate the convergence of the model. Most LLMs employ initialization techniques that were adopted in training smaller-scale models, such as conventional initialization techniques like (Kumar, 2017; He et al., 2015). For example, initialization method introduced by Kumar (2017) aims to balance input and output variances. Fixup (Zhang et al., 2019) and ZerO (Zhao et al., 2021) set the residual stem to zero, preserving signal identity. SkipInit (De and Smith, 2020) substitutes batch normalization with a zero-value multiplier. ReZero (Bachlechner et al., 2021) adds zero-valued parameters to maintain identity, leading to faster convergence. T-Fixup (Huang et al., 2020) follows Fixup to adopt rescaling schemes for the initialization of residual blocks of Transformer models. DeepNet (Wang et al., 2022c) adjusts the residual connection in deep Transformers using Post-LN-init, ensuring stable inputs to Layer-Normalization and mitigating gradient vanishing for stable optimization.

### v4

@@ 2.1.2 Parameter Pruning
Parameter pruning compresses LLMs by removing redundant or less important model weights. Parameter pruning methods for LLMs can be categorized into structured pruning and unstructured pruning.
@@
Structured Pruning. Structured pruning focuses on pruning structured patterns such as groups of consecutive parameters or hierarchical structures such as rows, columns, or sub-blocks of the LLM weight matrices. For instance, LLM-Pruner (Ma et al., 2023) introduces a task-agnostic structured pruning strategy that selectively eliminates non-essential interconnected structures using gradient information. LLM-Pruner utilizes a small amount of data to obtain the weight, parameter, and group importance of the coupled structure for LLaMA (Touvron et al., 2023a), and uses LoRA (Hu et al., 2022) to recover accuracy after pruning, showing competitive zero-shot performance. Sheared LLaMA (Xia et al., 2023), on the other hand, proposes two techniques to improve the performance of LLM-Pruner. The first technique, named targeted structured pruning, prunes a larger model to a designated target shape by eliminating layers, heads, intermediate and hidden dimensions in an end-to-end manner. The second technique, named dynamic batch loading, dynamically configures the composition of sampled data in each training batch based on losses in various domains. Through these two techniques, Sheared LLaMA is able to prune LLaMA2-7B down to 1.3B parameters, achieving superior compression ratio compared to LLM-Pruner. LoRAPrune (Zhang et al., 2023c) introduces a LoRA-based pruning criterion using LoRA’s weights and gradients for importance estimation. By employing an iterative structure pruning process to eliminate excess channels and heads, LoRAPrune achieves better efficiency over LLM-Pruner at 50% compression rate. Lastly, given the input, Deja Vu (Liu et al., 2023f) predicts a small set of attention heads and MLP parameters, referred to as contextual sparsity, yields approximately the same output as the dense model. By exploiting such contextual sparsity, Deja Vu is able to achieve much lower latency compared to FasterTransformer without accuracy drop.
@@
Unstructured Pruning. Unstructured pruning, on the other hand, focuses on pruning model weights individually. Compared to structured pruning, unstructured pruning has much more pruning flexibility and thus enjoys a lower accuracy drop. However, unstructured pruning incurs irregular sparsification, which in general makes the resulting pruned models difficult to be deployed on hardware except specific types of hardware such as Nvidia Ampere GPUs (Busato & Pool, 2020). For instance, Frantar & Alistarh (2023) introduce SparseGPT, an one-shot LLM unstructured pruning approach that does not require retraining. SparseGPT formulates pruning as a sparse regression problem and solves it by utilizing an approximate solver based on the inversion of the Hessian matrix. In doing so, SparseGPT reaches about 60 unstructured sparsity on models such as OPT-135B while experiencing only a slight performance drop. Sun et al. (2024) propose Wanda, which prunes weights based on the product values of weight magnitudes and their respective input activations. Compared to SparseGPT, Wanda neither relies on second-order information nor necessitates weight update, and is able to achieve competitive performance. Shao et al. (2024) improve the performance of SparseGPT in another way. Specifically, instead of performing the unstructured pruning with a unified ratio for every layer, they propose to utilize Hessian sensitivity-aware mixed sparsity pruning to achieve a minimum of 50% sparsity in LLMs without retraining. This method adaptively assigns sparsity based on sensitivity to minimize the error induced by pruning while preserving the overall level of sparsity.
2.1.3 Low-Rank Approximation
Low-rank approximation compresses LLMs by approximating the LLM weight matrix with smaller low-rank matrices and such that , where , , and is typically much smaller than . In doing so, low-rank approximation reduces the number of parameters and enhances efficiency. For example, Xu et al. (2023a) introduce TensorGPT which compresses the embedding layers of LLMs using Tensor-Train Decomposition (TTD). It transforms and breaks down each token embedding and creates an efficient embedding format named Matrix Product State (MPS) that can be efficiently computed in a distributed manner. LoSparse (Li et al., 2023g) improves the performance of TensorGPT by compressing the coherent and expressive components within neurons through low-rank approximation while eliminating the incoherent and non-expressive elements through pruning. As another line of research, FWSVD (Hsu et al., 2022) compresses the weight matrix of an LLM instead of the token embedding matrix via low-rank approximation. Specifically, instead of using vanilla singular value decomposition (SVD), FWSVD proposes a weighted SVD approach which uses Fisher information to weigh the importance of the weights for compression. While FWSVD demonstrates competitive compression results under low compression ratios, it requires calculating the gradients based on the training dataset of the target task to estimate the importance scores, which is task-specific and demands significant computation resources. In contrast, ASVD (Yuan et al., 2023b) proposes a training-free SVD-based approach. It scales the weight matrix based on the activation distribution that enhances the decomposition accuracy and efficiency for model compression. However, neither FWSVD nor ASVD directly correlate singular values with compression loss. Consequently, truncating the smaller singular values might result in increased compression loss. SVD-LLM (Wang et al., 2024c) addresses this drawback by incorporating a truncation-aware data whitening strategy that establishes a direct mapping between singular values and compression loss. Experimental results demonstrate the superiority of SVD-LLM over FWSVD and ASVD in terms of compression performance and speed.
2.1.4 Knowledge Distillation
Knowledge Distillation (KD) compresses LLMs by transferring knowledge from a large teacher LLM to a smaller student LLM. Though effective, compared to other LLM compression methods, knowledge distillation methods incur a resource-demanding distillation process. In general, KD for LLMs can be categorized into white-box KD methods and black-box KD methods.
@@ 2.1.4 Knowledge Distillation
White-Box Knowledge Distillation. White-box KD refers to KD techniques where the parameters or logits of the teacher LLM are used in the distillation process (Gou et al., 2021). For example, as a pioneering effort in this direction, Baby LLaMA (Timiryasov & Tastet, 2023) trains an ensemble of GPT-2 and a collection of smaller LLaMA models using the BabyLM dataset of 10M words. This ensemble is then distilled into a compact LLaMA model with 58 million parameters, which outperforms both its original teacher models as well as a comparable model that was trained without the use of distillation. Gu et al. (2024) observe that conventional KD objectives, such as Kullback-Leibler divergence (KLD), may not be suited for open text generation tasks due to their complex output spaces compared to classification tasks. To address this issue, they propose MiniLLM which minimizes reverse KLD using the gradient of the objective function through policy gradient techniques (Sutton et al., 1999). Experimental results show that MiniLLM achieves better accuracy than conventional KD or directly fine-tuning student models. KPTD (Padmanabhan et al., 2023) demonstrates that white-box KD can transfer and disseminate knowledge from entity definitions into the parameters of a pre-trained language model. Specifically, KPTD creates a transfer set by prompting the language model to generate text based on the definition of the entity. The model parameters are then updated to align the distribution of the student model with that of the teacher model. TED (Liang et al., 2023) introduces a technique for layer-specific task distillation. It uses specially designed filters to align the internal states of both student and teacher models in each layer. These filters extract relevant knowledge from the internal states that is beneficial for the specific task. TED shows considerable and steady gains in performance on both continual pre-training and fine-tuning. TSLD (Kim et al., 2023b) leverages token-level distillation to enhance QAT. It addresses the limitations of layer-to-layer KD in token prediction recovery by reforming intermediate representation and has successfully applied QAT to LLMs. MiniMA (Zhang et al., 2023a) proposes a viewport towards the capacity gap in distilling LLMs, converting it into a principle through analysis and introducing a 3B Language Model that sets a new benchmark for compute-performance pareto frontier. Experimental results show that MiniMA achieves the best accuracy compared to other 3B distilled LLMs. Lastly, Generalized knowledge distillation (GKD) (Agarwal et al., 2024) addresses the issue of distribution mismatch by drawing output sequences from the student model during training. GKD can be applied in combination with other distillation methods to improve their compression performance.
@@
| OPT (Zhang et al., 2022a) | 175B | 180B tokens | 992 A100-80G | - |
| BLOOM (Scao et al., 2023) | 176B | 366B tokens | 384 A100-80G | 105 days |
| GLM (Zeng et al., 2023) | 130B | 400B tokens | 786 A100-40G | 60 days |
@@
| Gopher (Rae et al., 2022) | 280B | 300B tokens | 1024 A100 | 13.4 days |
@@
| WeLM (Su et al., 2023a) | 10B | 300B tokens | 128 A100-40G | 24 days |
@@
Black-Box Knowledge Distillation. Different from white-box KD, in black-box KD, only the outputs generated from the teacher LLM are used in the distillation process. Inspired by ICT (Chen et al., 2022c) and MetaICL (Min et al., 2022a), where the language model is meta-trained under a wide range of tasks using in-context learning objectives and then fine-tuned for unseen tasks through in-context learning, Multitask-ICT (Huang et al., 2022) introduces a concept known as in-context learning distillation to transfer the few-shot learning capabilities from the teacher model to the student model. Experimental results show that under Multitask-ICT, in-context learning objectives achieve the best performance when combined with language modeling objectives. Similarly, Li et al. (2024b) introduce a hybrid prompting technique that employs multi-task learning along with explanations generated by GPT-3 text-davinci-002 version (OpenAI, 2023). This method is used to distill explanations into smaller models, achieving consistent and significant improvements over strong single-task fine-tuning benchmarks in different scenarios. Experiments on multiple reasoning tasks show that this method even perform better than finetuning or prompting a 60x larger GPT-3 (175B) model by up to 9.5% in accuracy. Lion (Jiang et al., 2023b) introduces an adversarial distillation architecture aimed at enhancing the efficiency of knowledge transfer by incrementally improving the skill level of the student model. Specifically, it prompts LLMs to recognize challenging instructions and creates new complex instructions for the student model, thereby establishing a three-phase adversarial cycle involving imitation, discrimination, and generation. Experimental results show that Lion-13B not only achieves comparable open-ended generation capabilities to ChatGPT but surpasses conventional instruction-tuned models. DISCO (Chen et al., 2023j) prompts a general LLM to produce phrasal perturbations. These generated perturbations are then filtered by a specialized teacher model to distill high-quality counterfactual data into smaller student models, allowing the smaller models to learn causal representations more reliably. As another line of research, some studies have shown that chain-of-thought (CoT) prompting can elicit language models to solve complex reasoning tasks step by step, with the aim to transfer such ability from large models into smaller ones through black-box KD. For example, to enhance the CoT math reasoning capabilities of smaller models, Fu et al. (2023b) propose a method for instruct-tuning a student model (FlanT5) by distilling the reasoning pathways found in the GSM8K dataset from a teacher model (GPT-3.5 code-davinci-002 (Chen et al., 2021b)). Fine-tuning and distilling smaller models require substantial amounts of training data to match the performance of the large model. To address this issue, Hsieh et al. (2023) propose Distilling Step-by-Step, a technique that uses CoT prompting to extract LLM rationales for extra guidance in training smaller models within a multi-task setting. Experimental results show that Distilling Step-by-Step achieves better performance with much fewer labeled or unlabeled training examples compared to both fine-tuning and standard distillation. Fine-tune-CoT (Ho et al., 2023) utilizes existing zero-shot CoT prompting techniques (Kojima et al., 2022) to create rationales from LLMs. These rationales are then used to fine-tune smaller student models. It also introduces diverse reasoning, a method that employs stochastic sampling to generate a variety of reasoning solutions from teacher models, which serves to enrich the training data for the student models. SOCRATIC CoT (Shridhar et al., 2023) breaks down the original problem into a series of smaller sub-problems and utilizes this decomposition to direct the intermediate steps of reasoning. It is used to train a pair of smaller, distilled models: one specializes in dissecting the problem and the other focuses on solving these sub-problems. SOCRATIC COT is shown to be an effective alternative to CoT, enabling a much smaller model (GPT-2 large) to outperform a 10x larger model (GPT-3 6B). SCOTT (Wang et al., 2023c) uses rationales generated by LLMs to train a student model under a counterfactual reasoning framework. It ensures that the student model does not overlook the provided rationales, thereby preventing it from making inconsistent predictions. Experimental results show that SCOTT can generate CoT rationales that are more faithful than original CoT prompting. Li et al. (2023b) present a method called symbolic CoT distillation (SCoTD) that draws CoT rationales from a LLM using unlabeled data instances. A smaller model is then trained to predict both the sampled rationales and the associated labels. Lastly, Peng et al. (2023a) utilize GPT-4 as a teacher model to generate English and Chinese instruction-based datasets to refine student LLMs. They show that the 52K data points generated by GPT-4 are able to improve zero-shot performance compared to instruction-following data generated from previous state-of-the-art models.
2.2 Efficient Pre-Training
As shown in Table 1, pre-training LLMs incurs significant costs. Efficient pre-training techniques focus on reducing the costs of the LLM pre-training process in terms of compute resources, training time, memory and energy consumption. As summarized in Figure 7, enhancing the efficiency of pre-training can be achieved through different and complementary techniques, including mixed precision acceleration, scaling models, initialization techniques, training optimizers, and system-level pre-training efficiency optimization.
@@ 2.2 Efficient Pre-Training
Mixed Precision Training, fill=blue!10
AMP (Micikevicius et al., 2018; Facebook AI Research (2023), FAIR; Rae et al., 2022),
Brain Floating Point (BFLOAT16) (Kalamkar et al., 2019; Burgess et al., 2019), leaf, text width=39em
@@
Progressive Stacking (Gong et al., 2019), MSLT (Yang et al., 2020), CompoundGrow (Gu et al., 2021), bert2BERT (Chen et al., 2022b),
Knowledge Inheritance (Qin et al., 2022),
Staged Training (Shen et al., 2022),
@@
Mango (Pan et al., 2023),
Yao et al. (2024),
Growth Strategy (Li et al., 2023e), leaf, text width=60em
@@
Kumar (2017), Fixup (Zhang et al., 2019), ZerO (Zhao et al., 2022), SkipInit (De & Smith, 2020),
ReZero (Bachlechner et al., 2021), T-Fixup (Huang et al., 2020), DeepNet (Wang et al., 2024a), leaf, text width=43em
Training Optimizers, fill=blue!10
Lion (Chen et al., 2023g),
Sophia (Liu et al., 2024a), leaf, text width=24em
@@
ZeRO-Infinity (Rajbhandari et al., 2021),
Zeus (You et al., 2023), Perseus (Chung et al., 2023), leaf, text width=60em
@@
Mixed Precision Training. Mixed precision training enhances pre-training efficiency by using low-precision models for forward and backward propagation and then converting the calculated low-precision gradients to high-precision ones for updating the original high-precision weights. For example, Micikevicius et al. (2018) propose Automatic Mixed Precision (AMP) to keep a master copy of weights in full-precision (FP32) for updates, whereas weights, activations, and gradients are stored in FP16 for arithmetic operations. Notably, the improved version of AMP (Facebook AI Research (2023), FAIR) has eliminated the copy of FP32 weights, but the optimizer (AdamW) still uses FP32 internally. Meanwhile, Rae et al. (2022) demonstrate that FP16 in AMP results in accuracy loss due to the restricted numerical range. To address this issue, Brain Floating Point (BFLOAT16), which has a greater dynamic range — i.e., number of exponent bits — than FP16, was proposed (Kalamkar et al., 2019; Burgess et al., 2019) to achieve better training performance.
@@
Scaling Models. Techniques based on scaling models accelerate pre-training convergence and reduce training costs by leveraging the weights of a smaller model to upscale to a larger one. For example, Gong et al. (2019) introduce a technique named progressive stacking to transfer knowledge from a simpler model to a more complex one to enhance model training efficiency. Meanwhile, Yang et al. (2020) observe that as the depth of the model increases through progressive stacking, the training speed however decreases. To address this issue, they propose multi-stage layer training (MSLT), which only updates the output and newly introduced top encoder layers while keeping the previously trained layers unchanged. Once all the layers have been trained, MSLT fine-tunes the entire model by updating each layer with 20% of the total steps, making it more time-efficient than the traditional progressive stacking approach. Similarly, Gu et al. (2021) introduce CompoundGrow, which begins with training a small model and then incrementally expands it using a mix of model growth techniques, including increasing input length, model breadth and depth, leading to an acceleration in the pre-training process in wall-clock time compared to progressive stacking. Chen et al. (2022b) propose bert2BERT, which applies function-preserving initialization (FPI) and advanced knowledge initialization (AKI) to transfer the knowledge of a smaller pre-trained model to a large model to improve the pre-training efficiency of the large model. Specifically, FPI enforces the initialized larger model to closely mirror the behavior of the smaller model, laying a strong basis for later optimization; and AKI promotes faster convergence by replicating weights from higher layers. Experimental results show that bert2BERT is able to save a significant amount of training cost over MSLT. Qin et al. (2022) propose Knowledge Inheritance which employs knowledge distillation as an auxiliary supervision during pre-training. This facilitates training a larger model from a smaller teacher model, thereby enhancing both the pre-training speed and the generalization ability. Shen et al. (2022) introduce Staged Training that begins with a small model and progressively increases its depth and breadth through a growth operator. By starting each stage with the results from the previous one, it effectively reuses computation, leading to a more efficient training process compared to previous techniques like CompoundGrow and progressive stacking. Wang et al. (2023d) propose Linear Growth Operator (LiGO) that linearly maps the parameters of a smaller model to initiate a larger one. By using a composition of width-and depth-growth operators further enhanced with Kronecker factorization to capture architectural knowledge, LiGO outperforms bert2BERT which saves about 30% computational costs. Pan et al. (2023) introduce a technique named Mango which establishes a linear relationship between each weight of the target model and all weights of the pretrained model to boost acceleration capabilities. It also employs multi-linear operators to decrease computational and spatial complexity during pre-training, achieving 59.9% acceleration ratio compared to Chen et al. (2022b) and LiGO. Drawing from these scaling techniques and the progressive pre-training (Yao et al., 2024), recent LLMs like FLM-101B (Li et al., 2023e) introduce a growth strategy to cut LLM training costs by expanding model structures offline and resuming from the previous stage’s smaller model checkpoint.
@@
Initialization Techniques. Initialization plays a key role in enhancing the efficiency of LLM pre-training because a good initialization can accelerate the convergence of the model. Most LLMs employ initialization techniques that were adopted in training smaller-scale models. For example, initialization method introduced by Kumar (2017) balances input and output variances. Fixup (Zhang et al., 2019) and ZerO (Zhao et al., 2022) set the backbone to zero, preserving signal identity. SkipInit (De & Smith, 2020) substitutes batch normalization with a zero-value multiplier. ReZero (Bachlechner et al., 2021) adds zero-valued parameters to maintain identity which leads to faster convergence. T-Fixup (Huang et al., 2020) follows Fixup to adopt rescaling schemes for the initialization of the residual blocks of Transformer models. DeepNet (Wang et al., 2024a) adjusts the residual connection in deep Transformers using Post-LN-init, ensuring stable inputs to layer normalization and mitigating gradient vanishing for stable optimization.